OpenStudy (rayman5766):

Lou makes pancakes at a restaurant. According to Lou’s recipe, the number of pancakes made is proportional to the number of eggs used. The table shows this relationship. What is the constant of proportionality? Eggs (x) 4 8 12 Pancakes (y) 18 36 54

OpenStudy (rayman5766):

is my answer 4.5

OpenStudy (igreen):

Yep. 18 / 4 = 4.5 36 / 8 = 4.5 54 / 12 = 4.5